OpenStudy (anonymous):

A ball is dropped from the top of a 1,000 foot building. The height of the ball is half its original height after each bounce. Part 1: What will the height of the ball be after 10 bounces? Part 2: Using complete sentences, explain the procedure taken to answer this question.

OpenStudy (anonymous):

okay.. this is interesting.. so 1,000/2 = 500 (1st bounce)

OpenStudy (anonymous):

then 500/2 = 250.. (2nd bounce)

OpenStudy (anonymous):

250/2 = 125 (3rd bounce)

OpenStudy (anonymous):

Can you do this part of the problem from here?

OpenStudy (anonymous):

Here you go, take a look at this :D http://answers.yahoo.com/question/index?qid=20110428204304AADqci8

OpenStudy (anonymous):

This seems to have what you're looking for.

OpenStudy (anonymous):

awe i was actually trying to do it lol but oh well thanxx

OpenStudy (anonymous):

It's ok! Now you know! THanks for the medal! Good luck in your studies!